# A Causal Framework to Measure and Mitigate Non-binary Treatment Discrimination

### v1

@@ 1 Introduction
loan repaid loan approved
@@
re-offence bail approved
@@
claim approval policy approved
@@
job performance job offered
Data-driven approaches increasingly leverage historical data to automate decision-making processes across societal domains, such as finance [4, 49], healthcare [20], and criminal justice [13]. These systems typically reduce complex decision-making to a binary classification task, predicting either ground-truth outcomes (e.g., loan repayment, recidivism after bail) or imperfect proxies (e.g., healthcare expenditures as a proxy for health status) [44, 53]. However, real-world decisions are rarely as simple as a binary choice; they involve multiple, non-binary treatment decisions. As shown in Table 1, decision-makers often determine not only whether to give a positive decision (provide treatment, e.g., approve loan) but also its specific terms (e.g., loan amount, duration, interest rate). Similar oversimplifications occur in other domains, such as bail decisions [61]. This simplistic approach overlooks the impact of non-binary treatment decisions on outcomes, leading to fairness analyses that fail to provide a holistic view of existing systems [46].
Despite growing research on fairness in algorithmic decision-making, existing analyses simplify complex decision-making mechanisms in three key ways. First, many (if not most) studies [69, 21, 11] include non-binary treatments as covariates of the decision-subject (covariate examples in Table 1). This assumption obscures the mechanisms that assign treatment decisions, which are not controlled by decision-subjects but by decision-makers and, thus, may reveal discriminatory patterns. This assumption thus limits the ability of these works to account for all sources of unfairness and their implications. Second, alternative works [39, 12, 57] examine the fairness implications of treatment decisions but assume they are binary, ignoring the multivariate, non-binary nature of treatment decisions (such as, e.g., loan or bail amount). Finally, only a few works [2, 17, 3] have focused on disparities in non-binary treatment decisions across different groups. Still, they lack a generalizable framework, providing only empirical insights in financial domains and, more importantly, neglecting the discriminatory downstream effects of these disparities on outcomes.
To bridge gaps in the existing approaches, we propose a fundamental paradigm shift in fairness analysis of automated decision-making systems that brings fairness considerations to non-binary treatment decisions and their downstream outcome effects. Our paradigm shift leverages the following key insight: decision-makers do not simply make binary decisions (whether to give treatment) but also control the assignment of adjacent non-binary treatment decisions, which can introduce disparities and negatively impact outcomes. To incorporate this insight into fairness analysis, we introduce a broadly applicable causal framework that allows decision-makers to analyze the fairness of underlying decision-making mechanisms systematically. This systematic analysis allows our framework to provide novel measures of treatment disparities, assess their discriminatory effects on outcomes, and enable mitigation strategies to prevent biases in non-binary treatment decisions from perpetuating unfairness in automated systems.
Our approach explicitly differentiates between a decision-subject’s covariates and a decision-maker’s treatment decisions in the causal data-generating process, as illustrated in Figure 1. This causal formulation allows us to measure disparities in non-binary treatment decisions $Z$ induced by sensitive attributes $S$, both directly (solid red path) and indirectly through covariates $X$ (dotted red path). Furthermore, our framework quantifies how these disparities propagate to outcomes $Y$ via the causal path from $Z$ to $Y$ (dashed orange path), thereby inducing treatment discrimination. Our framework supports mitigation in scenarios impacted by treatment discrimination through two pre-processing strategies. These techniques generate treatment-fair counterfactual datasets, enabling fairer non-binary automated decisions and more equitable risk score estimation to inform future decision-making.
Our framework advances fairness analyses by leveraging path-specific counterfactuals to quantify direct and indirect disparities in non-binary treatment decisions and their impact on outcomes. While path-specific counterfactuals provide a powerful tool for causal fairness analyses, existing methods [52, 8] suffer from significant limitations, such as being restricted to linear data generation processes or requiring inefficient Monte Carlo estimation with strong regularization. To overcome these challenges, we build on the state-of-the-art in causal inference, the causal normalizing flows [26], and propose a novel path-specific counterfactual generation method. This advancement enables our framework to be theoretically grounded while being practical to be applied in real-world decision-making systems.
Finally, we validate our framework using four real-world lending datasets across two decision-making scenarios. Our empirical results demonstrate that explicitly accounting for non-binary treatment decisions helps provide more holistic fairness assessments, identifies systemic biases, and supports mitigation efforts. These findings underscore the importance of addressing treatment discrimination in achieving fairer outcomes, bridging a critical gap in fairness research, and laying the foundation for more robust causal fairness analyses.
$X$$Z$$Y$$S$
Figure 1: Causal graph of the data generation process, distinguishing decision-subject covariates $X$ and treatment decisions $Z$. Under the assumption that a positive binary decision has already been made. Red paths (solid: direct, dotted: indirect) show causal paths to $Z$ influenced by sensitive attribute $S$, highlighting potential sources of disparities. The orange dashed path from $Z$ to $Y$ captures the potentially discriminatory effects on the outcome.
2 Background
Our framework for measuring and mitigating treatment discrimination in decision-making relies on causal analyses. This section briefly outlines the core causal concepts used throughout the paper.
Structural causal models (SCMs). An SCM [56] $\mathcal{M}=(V,U,F)$ is defined by the observational features $V$, independent, unobserved, exogenous variables $U$ (no hidden confounding) and structural equations $F$. A structural equation for some $V_{i}$ gives the functional mapping from the corresponding exogenous $U_{i}$ and causal parents $\operatorname{pa}(V_{i})$ to the value of $V_{i}$ as $V_{i}=F_{i}(\operatorname{pa}(V_{i}),U_{i})$. These cause-effect relations in $\mathcal{M}$ can also be represented by a directed acyclic graph $\mathcal{G}$ like in Figure 1, where the nodes represent $V$, and the directed edges represent the cause-effect relations.
Performing interventions. The SCM $\mathcal{M}$ helps us estimate the effects of hypothetical interventions $\mathcal{I}$ on features. For instance, to analyze the effect of the intervention $\mathcal{I}$ setting $V_{i}=\alpha$, we use the ${\operatorname{do}}$-operator [56] ${\operatorname{do}}(V_{i}\rightarrow\alpha)$ on $\mathcal{M}$ that generates the intervened SCM $\mathcal{M}^{\mathcal{I}}$.
Counterfactuals. In the absence of hidden confounders, this intervened SCM $\mathcal{M}^{\mathcal{I}}$ can also be used to perform counterfactual analyses for a particular factual decision-subject with features $\vartheta^{F}$ to ask “what would be the counterfactual features $\vartheta^{CF}$ had we intervened with ${\operatorname{do}}(V_{i}\rightarrow\alpha)$, and all exogenous factors $u^{F}$ remaining unchanged.” Counterfactuals $\vartheta^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha))$ are computed following the three steps [56]: i). abduction: estimating $u^{F}$ from $\vartheta^{F}$ using $\mathcal{M}$, ii). action: performing ${\operatorname{do}}(V_{i}\rightarrow\alpha)$ to get $\mathcal{M}^{\mathcal{I}}$, iii). prediction: compute $\vartheta^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha))$ from $\mathcal{M}^{\mathcal{I}}$ using $u^{F}$ and setting $V_{i}=\alpha$. In simple terms, computing counterfactuals estimates the total effect of an intervention on the downstream causal features of the intervened variable in the causal graph $\mathcal{G}$ for the factual decision-subject $\vartheta^{F}$.
Path-specific counterfactuals. In certain scenarios, it is crucial to understand the effects of a particular feature $V_{i}$ on a downstream feature $V_{j}$ along different causal paths in $\mathcal{G}$. This analysis can be done by computing path-specific counterfactuals [8], which use multiple ${\operatorname{do}}$ operations to control for specific causal paths. Consider the two paths in an example causal structuring between variables $\langle V_{i},V_{j},V_{k}\rangle$, where we have $V_{i}\rightarrow V_{k}$ and $V_{i}\rightarrow V_{j}\rightarrow V_{k}$. For a factual instance $\vartheta^{F}=\langle v_{i}^{F},v_{j}^{F},v_{k}^{F}\rangle$, we can compute the path-specific counterfactual effect of ${\operatorname{do}}(V_{i}\rightarrow\alpha)$ along these paths. For the path $V_{i}\rightarrow V_{k}$, we measure the counterfactual effect using $v_{k}^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha),{\operatorname{do}}(V_{j}\rightarrow v_{j}^{F}))$. Similarly, for the path $V_{i}\rightarrow V_{j}\rightarrow V_{k}$, we use $v_{k}^{CF}(v_{i}^{F},{\operatorname{do}}(V_{j}\rightarrow v_{j}^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha))))$. For the path $V_{i}\rightarrow V_{j}\rightarrow V_{k}$, we use the counterfactual $v_{j}^{CF}$ obtained from the intervention on $V_{i}$ but keep $V_{i}=v_{i}^{F}$ fixed along the other path.
Causal fairness analyses. The SCM framework has been extensively used to causally study fairness problems in ML [35, 8]. In these scenarios, target labels $Y$ (e.g., bail, loan repayment) are considered causally affected by the sensitive feature $S$ (e.g., gender/sex, race). We can estimate the total effect of $S$ on $Y$ by finding sensitive counterfactuals $y^{\mathrm{SCF}}=y^{CF}({\operatorname{do}}(S\rightarrow\alpha))$, as was shown in counterfactual fairness [35]. Similarly, Chiappa [8] used path-specific counterfactuals $y^{\mathrm{PSCF}}$ to analyze the potentially discriminatory effect of $S$ on $Y$ along different causal paths in $\mathcal{G}$ for linear SCMs. In this work, we extend the concept of counterfactual fairness and path-specific counterfactual fairness for treatment decisions and their downstream effects.
3 Measuring treatment disparities and their downstream effects
In this section, we introduce our causal framework to address the limitations of the existing fair decision-making literature by explicitly measuring disparities in treatment decisions and their downstream effects and mitigating potential discriminatory consequences. To this end, we assume that the data generation process that produces the observed (factual) historical data $\mathcal{D}$ follows the causal graph depicted in Figure 1, where:
@@ 3 Measuring treatment disparities and their downstream effects
    The treatment decision $Z$ is a causal child of both $X$ and (potentially) $S$. We assume that the causal mechanism that generated $Z$ is under the control of the decision-maker (e.g., we assume that banks decide not only whether or not to grant a loan but also the loan terms) and may be discriminatory. The direct causal path from $S$ to $Z$ is shown in solid red, while the indirect path mediated by $X$ is shown in dashed red.
@@
    The treatment decision $Z$ is a causal parent of the outcome variable $Y$ (e.g., we assume loan terms can impact the repayment ability of applicants). Hence, $S$ can influence $Y$ mediated by $Z$ (orange dashed line in Figure 1), leading to the possible propagation of treatment disparities to the outcome. Moreover, we maintain the direct $S\rightarrow Y$ since such effects may occur in specific scenarios, e.g., in healthcare.
Remarks. The edges of our causal graph $\mathcal{G}$ (Figure 1) indicate the structural causal ordering [26] among the different groups of features, where the cause-effect relation for some edges may be zero. Moreover, although the figure shows a single sensitive feature for simplicity, our $\mathcal{G}$ can consider multiple sensitive features as independent root nodes. We treat $X$ and $Z$ in $\mathcal{G}$ as multivariate blocks, focusing our causal analyses on relations between these “blocks” without making explicit assumptions about their causal relations.
3.1 Disparities in non-binary treatment decisions
Given our proposed causal graph in Figure 1, we use the SCM framework to measure disparities in treatment decisions by measuring counterfactual and path-specific counterfactual effects of the sensitive feature (e.g., gender) on the treatment decision variables. For simplicity and without loss of generalization, here we assume the sensitive feature on which we intervene to be binary, i.e., $S\in\{0,1\}$.  [^1] Our measures can easily be extended for multi-valued categorical sensitive attributes, e.g., by conducting group-pairwise comparisons. Intersectional disparity measurements across multiple sensitive attributes are left for future work. Given an observed dataset $\mathcal{D}=\{s^{F},x^{F},z^{F},y^{F}\}$, disparity function $\mathcal{F}$, and aggregation function $\mathcal{A}_{\mathcal{D}}$ regarding $\mathcal{D}$, we define two treatment disparity measures.
Definition 1.
Total Treatment Disparity (TTD), against decision-subjects with (factual) sensitive feature $s^{F}$, measures the expected difference between the treatment $z^{F}$ they received and the treatment $z^{\mathrm{SCF}}$ they would have received if they belonged to a different sensitive group, i.e.,
$\operatorname{TTD}(s^{F})=\mathcal{A}_{\mathcal{D}}\left(\mathcal{F}(z^{\mathrm{SCF}},z^{F})\right),\,\mathrm{where}\,z^{\mathrm{SCF}}=z^{CF}({\operatorname{do}}(S\rightarrow(1-s^{F})))$ (1)
Related to [35], the ${\operatorname{do}}$-operator on the sensitive feature $S$ gives the sensitive counterfactual treatment $z^{\mathrm{SCF}}$ a decision-subject would have received if their sensitive group were different.
Definition 2.
Direct Treatment Disparity (DTD), against decision-subjects with (factual) sensitive feature $s^{F}$, measures the expected difference between the treatment they received and the sensitive (direct) - path counterfactual treatment, which assumes a ${\operatorname{do}}(S\rightarrow(1-s^{F})))$ only on the direct path from $S$ to $Z$, i.e.,
$\begin{split}\operatorname{DTD}(s^{F})=\mathcal{A}_{\mathcal{D}}\left(\mathcal{F}(z^{\mathrm{SPCF}},z^{F})\right),\,\mathrm{where}\,z^{\mathrm{SPCF}}=z^{CF}({\operatorname{do}}(S\rightarrow(1-s^{F})),{\operatorname{do}}(X\rightarrow x^{F}))\end{split}$ (2)
Leveraging path-specific counterfactuals [52, 8], we compute the sensitive path counterfactual treatments $z^{\mathrm{SPCF}}$ a decision-subject would have received for different sensitive group memberships while keeping original covariates $x^{F}$ fixed.
Remarks. First, the nature of $\mathcal{A},\mathcal{F}$ in the above definitions depends on the nature of the treatment decision variables. For instance, when treatment decisions are continuous, $\mathcal{A}$ would be the expectation $\mathbb{E}$ over $\mathcal{D}$, and $\mathcal{F}$ would denote the differences across counterfactual treatments. Second, by drawing parallels with predictive fairness concepts, we can better understand the fairness implications of treatment disparities. For instance, $\operatorname{DTD}$ highlights disparities caused by directly incorporating $S$ into treatments, which may be acceptable in areas such as healthcare but undesirable in contexts such as lending, aligning with the predictive fairness concept of disparate treatment [15]. Similarly, zero $\operatorname{TTD}$ requires that treatments $Z$ be independent of $S$, showing parallels with demographic parity [15], which demands that predictions do not depend on $S$. However, like predictive fairness, the relevance of perfect treatment parity depends on context. Similar to demographic parity scenarios, when ground-truth outcome $Y$ is unavailable (e.g., admissions, hiring), treatment disparities may be considered discriminatory. On the other hand, as separation-based frameworks [21, 69] show, the outcome $Y$ can explain predictive disparities in certain contexts, e.g., when $Y$ is a ground truth as in lending. Hence, studying the downstream impact on $Y$ is imperative to understand if treatment disparities are discriminatory.
3.2 Downstream effects of treatment disparities
While Definitions 1 and 2 help measure treatment disparities, it is crucial to assess the downstream impacts of these disparities on the outcome $Y$ (e.g., when $Y$ is a ground truth as in lending) to determine whether they are discriminatory. Hence, we introduce two novel measures to quantify the counterfactual downstream outcome effects of total and direct treatment disparities.
Definition 3.
where $\mathbb{I}$ is an indicator function. This definition helps us understand how the total causal effect of $S$ on the treatment decisions $Z$ ends up impacting the decision outcomes $Y$. Intuitively, for any decision-subject belonging to $s^{F}$ with covariates $x^{F}$, the decision-maker provides the sensitive counterfactual $z^{\mathrm{SCF}}$ treatments to measure the outcome change.
Definition 4.
$\operatorname{DTD-E}(s^{F},y^{F})$ $=\mathbb{E}_{\mathcal{D}}\left[\mathbb{I}\left(y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SPCF}}))\neq y^{F}\right)\right],$ (4)
where $\mathbb{I}$ is an indicator function. This definition helps us understand how the direct causal effect of $S$ on the treatment decisions $Z$ ends up impacting the decision outcomes $Y$. Intuitively, for any decision-subject belonging to $s^{F}$ with covariates $x^{F}$, the decision-maker provides the sensitive path counterfactual $z^{\mathrm{SPCF}}$ treatments to measure the outcome change.
Remarks. As before, we establish connections to predictive fairness concepts to understand when treatment disparities can result in discrimination through unfair decision outcomes. For specific contexts, we can relate to the perspective of disparate treatment [15] and consider any direct use of $S$ in assigning $Z$ problematic and any consequent effect on $Y$ as discriminatory. In other scenarios, e.g., healthcare, such effects may not be discriminatory if the effect of DTD on $Y$ is beneficial. Similarly, disparities between treatment decisions $z^{\mathrm{SCF}}$ and $z^{F}$ can lead to unfairness in $Y$. For instance, in predictive fairness, separation [69] deems decision disparities justifiable if they do not result in adverse ground-truth outcomes. Similarly, treatment disparities regarding $S$ may be deemed discriminatory if they do not have a positive impact on the ground-truth $Y$, e.g., when assigning $z^{F}$ results in negative outcomes ($y^{F}$), whereas $z^{\mathrm{SCF}}$ would have led to more favorable outcomes ($y^{CF}$).
3.3 Practical implementation
Prior works that computed path-specific counterfactuals for predictive fairness [52, 8] are restricted to linear settings or must rely on imprecise Monte Carlo approximations and complex regularization. To compute path-specific counterfactuals in more practical settings, we propose a novel framework that helps estimate our proposed disparity measures. Assuming access only to factual (potentially unfair) data $\mathcal{D}=\{s^{F},x^{F},z^{F},y^{F}\}$ and the causal graph $\mathcal{G}$ (Figure 1), our framework builds on the causal normalizing flows (CNF) [26]. CNF has demonstrated accurate causal estimates in societal decision-making settings [41]. Hence, by building our framework based on CNF, we develop a practical yet theoretically grounded approach to compute path-specific counterfactuals.
3.3.1 Background on causal normalizing flows
Causal normalizing flows (CNF) approximate the true and unknown SCM $\mathcal{M}$ using a single, invertible normalizing flows model $T_{\psi}$ with parameters $\psi$. Note that CNF can handle partial causal graphs, in which each random variable in the graph may be multivariate without explicitly modeling the causal relationships across its dimensions. Moreover, given a causal ordering across the random variables, CNF is highly flexible in learning the presence or absence of causal effects from observational data. Hence, even with data collected from a (potentially unfair) decision-making process, the CNF model provides an ideal practical framework to estimate our discrimination measures regarding treatment decisions and their downstream effects. Importantly, a causal normalizing flows model fitted on observational data can be used to estimate the counterfactual datapoint $\vartheta^{CF}$ resulting from a causal intervention $({\operatorname{do}}(V_{j}\rightarrow\alpha))$ on a given factual datapoint $\vartheta^{F}=\{v_{1}^{F},\ldots,v_{K}^{F}\}$ via:
  1. 1.
    abduction step computing the exogenous values for $\vartheta^{F}$ as ${u}^{F}=T_{\psi}(\vartheta^{F})$;
  2. 2.
    action step constructing the exogenous values for $\vartheta^{CF}$ as ${u}^{CF}({\operatorname{do}}(V_{j}\rightarrow\alpha))=\{{u}^{F}_{1:j-1},T_{\psi}({\operatorname{do}}(V_{j}\rightarrow\alpha)),{u}^{F}_{j+1:K}\}$;

### v2

@@ B.1 Treatment Decisions Across Different Domains
loan approved loan repaid
@@
bail approved re-offence
@@
policy approved claim approval
@@
job offered job performance
Table 4 details treatment decisions and covariates across various domains relevant to algorithmic fairness. For each domain, we identify the treatment decision features, which are those that are under the decision-maker’s control and directly influence the outcome. The table also specifies the binary decision for each domain. Importantly, these binary decisions may serve as either ground truth labels or proxies for the underlying outcome of interest. For example, the criminal justice domain: while automated systems typically predict bail approval decisions, the true objective is assessing defendant risk. Judges use this risk assessment to make the binary bail decision and consequently determine bail terms. However, focusing solely on the fairness of the binary decisions of bail approval might fail to address the unfairness in the bail terms.
Existing research in the social sciences and medicine has shown potential biases in treatment decisions. For instance, legal scholars have shown potential bias issues in bail treatment decisions (Arnold, Dobbie, and Yang 2018), while some quantitative research has shown unexplained gender biases in insurance costs (Pernagallo, Punzo, and Torrisi 2024). Similarly, Castilla (2008) empirically analyzed potential gender disparities in salary hikes in organizations. In healthcare, potential biases were also shown to exist in several treatment factors (Hershman et al. 2005).
Our proposed causal framework can be readily used to study treatment disparities and outcome effects across all these representative decision-making domains. Importantly, our framework can allow for a more comprehensive evaluation of such real-world scenarios using state-of-the-art computational methods. However, accessing appropriate datasets across most domains remains a significant challenge, as stated by existing research (Pernagallo, Punzo, and Torrisi 2024). Even when some data is available, they lack sufficient fine-grained information about treatment decisions provided by the decision-makers. With future advancements in data availability, our proposed causal framework can be used to study potential treatment disparities and fairness issues across multiple domains.
B.2 Background on Causal Normalizing Flows
Causal normalizing flows (CNF) (Javaloy, Sánchez-Martín, and Valera 2024) enable causal inference from observational data under minor assumptions. It provides an ideal practical methodology to estimate unknown causal relations from observational data and an accurate technique to estimate interventional and counterfactual quantities. Here, we provide a brief background on CNFs and refer the reader to the original paper for full details.
Assumptions
The work assumes the following for its theoretical grounding. First, it assumes a diffeomorphic data generation process. Hence, the structural equation functions $F$ in the SCM are invertible, and the functions and their inverse are differentiable. Second, it assumes the causal mechanism has no feedback loops; hence, it can be represented by a directed acyclic graph (DAG). Third, it assumes causal sufficiency. Hence, the exogenous variables for each feature are mutually independent, i.e., $p(U)=\prod_{i}p(u_{i})$. Finally, it assumes that a causal ordering is given across the features.
Autoregressive architecture
Under the assumptions, the authors show that SCM mechanisms correspond to a special class of autoregressive functions known as triangular monotone increasing (TMI) maps. Through TMI maps, the SCM is represented by triangular, invertible functions. Each feature in $V$ is a monotone function of its parents $pa(V)$ and its exogenous noise term $u$. These maps can be estimated using autoregressive normalizing flows (ANFs), where each conditional is modeled as a monotone transformation, forming the causal normalizing flows model.
The CNF defines a parameterized function $T_{\psi}$ that maps exogenous variables $U$ to observed features $V$, serving as a generative model for the features. Leveraging the autoregressive structure, the inverse $T_{\psi}^{-1}$ maps $V$ back to the exogenous space $U$. Since we assume the causal topological order is known, each conditional in the flow corresponds directly to a structural equation in the SCM.
By modeling the SCM in this way, CNFs can flexibly approximate the joint distribution over observed variables while remaining consistent with the underlying causal structure. For further details on the architecture and training procedure, we refer to the main paper (Javaloy, Sánchez-Martín, and Valera 2024).
Discussion on Theoretical Guarantees
Under the stated assumptions and leveraging the flow-based approximation of the SCM, the authors show that if the CNF $(T_{\psi},p(U))$ reproduces the observational distribution generated by the true SCM, then the exogenous variables learned by the flow differ from the true exogenous noise variables only by component-wise invertible transformations. Specifically, for each feature $i$, the corresponding exogenous factor learned by the CNF is a transformation $h_{i}(u_{i})$ of the true exogenous variable $u_{i}$, where $h_{i}$ is an invertible function. This result establishes that CNFs guarantee identifiability of exogenous variables up to invertible transformations, and by successfully isolating the exogenous information, they ensure causal consistency in the learned generative process.
Interventions and do-operation
While the traditional three-step process of Pearl (Pearl et al. 2000) manipulates the features directly for interventions, CNFs perform interventions by manipulating the latent exogenous variables. The process is discussed in Section 3.3. For generating counterfactuals, we first use the flows model to map the observed features $\vartheta^{F}$ to exogenous factors as $u=T_{\psi}(\vartheta^{F})$. Next, for the intervention ${\operatorname{do}}(V_{j}\rightarrow\alpha)$, we recompute the exogenous factor specifically for feature $j$, hence, $u_{j}=T_{\psi}({\operatorname{do}}(V_{j}\rightarrow\alpha))$. Hence, in the exogenous space, only for the exact intervened feature, the exogenous variable is updated. For all other features, they remain the same. Finally, one uses the inverse of the flows model $T^{-1}_{\psi}$ to go from post-interventional $u$ to $\vartheta^{CF}$. Further details on the $\operatorname{do}$}-operation in the CNF can be found in Appendix C in (Javaloy, Sánchez-Martín, and Valera 2024).
Handling partial causal graphs
A key strength of CNFs, which we leverage in our work, is their ability to learn from partial causal graphs, where the complete causal graph is unknown, but the causal ordering between subsets of observed features is available. Each subset of features with unknown internal ordering can be treated as a block, referred to in (Javaloy, Sánchez-Martín, and Valera 2024) as a strongly connected component (SCC).
Given knowledge of the causal ordering between these blocks, the CNF applies autoregressive flow modeling across the blocks. Within each block, the joint distribution is modeled without assuming any specific ordering. Consequently, the CNF retains its theoretical guarantees and identifiability properties at the block level, while identifiability is not ensured within individual blocks.
In our causal framework, we adopt this formulation by treating all covariates $X$ and all treatment decisions $Z$ as separate blocks. Through our counterfactual analysis of treatment and outcome disparities, we evaluate block-level counterfactuals for covariates and treatments to study their causal effects on the downstream outcome $Y$.
B.3 Extending to Multiple Sensitive Attributes
Our definitions for measuring treatment disparities and their outcome effects (Definitions 1-4) can be naturally extended to scenarios where we have multi-valued and multiple sensitive attributes, e.g, non-binary gender identities, race/ethnicity, religion, etc.
Let us consider $K$ different sensitive attributes, where each attribute $S_{i}$ has unique values $\mathcal{S}_{i}$. Note that this naturally considers binary and multi-valued categorical sensitive attributes. The combined set of all possible sensitive attribute values can be denoted as $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{2}\times\dots\times\mathcal{S}_{K}$. To compute counterfactual treatment decisions and outcomes, a central challenge is interpreting what constitutes a meaningful counterfactual across multiple sensitive attributes. Multiple plausible counterfactuals may exist (depending on which attributes are considered), and we outline two different strategies here.
  • Single flip. We change one sensitive attribute at a time (e.g., only gender), holding the others fixed to their factual values. For example, what treatment would an individual receive if their gender were different but their race and religion stayed the same? For attribute $S_{i}$, the counterfactual treatment can be defined for any value $s^{\prime}_{i}$ as:
    $z^{\mathrm{SCF}}=z^{CF}({\operatorname{do}}(S_{i}\rightarrow s^{\prime}_{i})),\forall\,s^{\prime}_{i}\in\mathcal{S}_{i}\setminus\{s^{F}_{i}\}$
  • Joint flip. We simultaneously change all sensitive attributes to a new disadvantaged group (changing the value set of $S$ from $s^{F}$ to a different value set $s^{\prime}$). For example, an individual who is factually a White male Christian would be compared to a Black female Muslim. Formally, the counterfactual treatment would be:
    $z^{\mathrm{SCF}}=z^{CF}({\operatorname{do}}(S\rightarrow s^{\prime})),\forall\,s^{\prime}\in\mathcal{S}\setminus\{s^{F}\}$
Our framework accommodates both counterfactual strategies, enabling interventions ${\operatorname{do}}(S\rightarrow s^{\prime})$ across multiple sensitive attributes using the estimated CNF model. However, each choice corresponds to a distinct fairness question of interest, whether one is concerned with the effects of individual attributes or disparities across entire groups. The work of (Wastvedt, Huling, and Wolfson 2024; Alvarez and Ruggieri 2025) defines counterfactuals involving multiple sensitive attributes through a joint flip, whereas (Yang, Loftus, and Stoyanovich 2020) defines counterfactuals relative to a chosen reference subgroup.
Treatment disparities.
Since we now need to consider potentially multiple sensitive counterfactuals across different sensitive attributes and values, we have to extend our definitions to aggregate disparities properly. For treatment-level disparities (Definitions 1 and 2), we can apply an aggregation operator $\mathcal{A}$ over the absolute treatment differences $\big|\Delta(z^{\mathrm{SCF}}_{s^{\prime}},z^{F})\big|$ before taking the expectation over the dataset $\mathcal{D}$. For instance, the multi-valued TTD would be:
$\operatorname{TTD}_{\text{M}}(s^{F})=\mathbb{E}_{\mathcal{D}}\Big[\mathcal{A}\big(\{|\Delta(z^{\mathrm{SCF}}_{s^{\prime}},z^{F})|\}_{s^{\prime}\in\mathcal{S}\setminus\{s^{F}\}}\big)\Big].$ (9)
Following (Wastvedt, Huling, and Wolfson 2024), we can consider different types of aggregate functions.
@@ B.3 Extending to Multiple Sensitive Attributes
    Average:
    $\mathcal{A}_{\text{AVG}}(s^{F})=\frac{1}{(|\mathcal{S}|-1)|\mathcal{S}|/2}\sum_{s^{\prime}\in\mathcal{S}\setminus\{s^{F}\}}\big|\Delta(z^{\mathrm{SCF}}_{s^{\prime}},z^{F})\big|$
@@
    Maximum absolute error:
    $\mathcal{A}_{\text{MAX}}(s^{F})=\max_{s^{\prime}\in\mathcal{S}\setminus\{s^{F}\}}\big|\Delta(z^{\mathrm{SCF}}_{s^{\prime}},z^{F})\big|$